# A method to detect the VUV photons from cooled 229Th:CaF2 crystals

### v1

@@ Abstract
Thorium-229, with its exceptionally low-energy nuclear excited state, is a key candidate for developing nuclear clocks. 229Th-doped CaF2 crystals, benefiting from calcium fluoride’s wide band gap, show great promise as solid-state nuclear clock materials. These crystals are excited by vacuum ultraviolet (VUV) lasers, which over time cause radiation damage. Cooling the crystals can mitigate this damage but introduces a challenge: photoabsorption. This occurs when residual gas molecules condense on the crystal surface, absorbing VUV photons and deteriorating detection efficiency. To solve this, we developed a cooling technique using a copper shield to surround the crystal, acting as a cold trap. This prevents ice-layer formation, even at temperatures below $-100\,^{\circ}$, preserving high VUV photon detection efficiency. Our study detailed the experimental cooling setup and demonstrated the effectiveness of the copper shield in maintaining crystal performance, a critical improvement for future solid-state nuclear clocks operating at cryogenic temperatures.
@@ 2.1 VUV photon detection setup
Our experimental setup [15, 18] was developed to produce the 229Th isomer states by using 29 keV X-ray and detect the signal photons, at the time when direct spectroscopy by VUV laser was unpromising [19, 20], due to the large energy uncertainty of the isomer state at that time. This report will introduce the updated setup for cooling the target crystal.
Figure 2: (a) The radioluminescence spectrum of the 229Th:CaF2 crystal measured by TU Wien group [21]. (b) Efficiencies of optical devices used in this measurement. Q.E.: quantum efficiency. Q.E. data were taken from Hamamatsu Photonics K. K. [22]. DCM reflectance is the total reflectance of 4 dichroic mirrors, measured by our group.
@@
To study the ice-layer formation and its impacts on isomer detection, we use the RL from the target crystal to assess the signal detection efficiency of our setup. As shown in Fig. 2 (a), the Cherenkov radiation includes a VUV component, which closely resembles the isomer signal [21, 25, 26]. This VUV component can be selected using the preferential reflective band of the DCM assembly, as shown by the red curve in Fig. 2 (b). Since both the RL and isomer photons are emitted within the crystal, we use the selected Cherenkov photons as a VUV source to evaluate our setup’s efficiency for detecting isomer signals.
@@ 2.2 Cooling setup
The purpose of the cooling experiment is to lower the crystal’s temperature to approximately 100 K, reducing the damage by laser [12] or X-ray irradiation. Directly attaching the small crystal to the cold copper head or using a copper holder can achieve this target temperature; however, such methods are not feasible. Because we need to frequently install/uninstall, and anneal (400 C) [18] the target crystal, and we need to change the position of the crystal to gain maximum detection efficiency. Instead, we cool the stainless steel target holder to achieve the cooling of the crystal. As illustrated in Fig. 1 (c), the crystal ($\sim$1 mm3) is mounted on the front surface of the target holder, which has a hole in its body. A flexible copper wire connects the target holder to a copper rod or bar, enabling positional adjustments. The opposite end of the copper rod or bar is fixed to the cold head. The insulator and 3D stage arm are made of PEEK plastic to mitigate heat conduction from the stage.
@@ 2.3 Vacuum improvement practices
When we started the cooling experiment, the vacuum condition was 10-2 Pa, and at that time the detection rate of the VUV-PMT decreased very rapidly in the initial cooling test, due to the ice-layer absorption. This relatively low vacuum condition was due to several factors: the low pumping speed of the evacuation connection, leakage of the flanges, and contamination-induced outgassing and evaporation within the chamber. This vacuum condition impedes the cooling experiments essential for the studies of 229Th isomer, necessitating efforts to address this issue.
@@ 2.4 Improved cooling scheme: copper rod conductor and shield
The shield is designed to absorb residual molecules and prevent them from condensing on the crystal surface. As shown in Fig. 1 (b), one end of the shield is closed off with the parabolic mirror with 1 mm gap kept between the mirror and the copper shield to avoid contact when the copper shrinks at low temperatures. The other end is completed with the MgF2 window which separates the vacuum between the irradiation chamber and the detection chamber. The parabolic mirror, copper shield, and MgF2 window together enclose the target holder, creating a semi-closed space free from molecular flow. Meanwhile, the cold shield works as a cold trap that captures the molecules inside it, therefore it features high vacuum conditions for the crystal cooling. The vacuum level reaches $5\times 10^{-6}$ Pa during operation of the cooler.
@@ 3.1 Cooling test
As seen in panel (b), the introduction of the copper shield and the MgF2 window significantly suppresses the ice-layer formation, as indicated by the stable VUV-PMT counting rate when the crystal is cooled below $-100\,^{\circ}$. Therefore, during the full cooling and warming cycle, the VUV-PMT count rate was unaffected by layer absorption.

### v2

@@ Abstract
Thorium-229, with its exceptionally low-energy nuclear excited state, is the only candidate for developing a nuclear clock. 229Th-doped CaF2 crystals, benefiting from calcium fluoride’s wide band gap, show great promise as solid-state nuclear clock materials. These crystals are excited by vacuum ultraviolet (VUV) lasers, which over time cause radiation damage. Cooling the crystals can mitigate this damage but introduces a challenge: photoabsorption. This occurs when residual gas molecules condense on the crystal surface, absorbing VUV photons and deteriorating detection efficiency. To solve this, we developed a cooling technique using a copper shield to surround the crystal, acting as a cold trap. This prevents ice-layer formation, even at temperatures below $-100\,^{\circ}$, preserving high VUV photon detection efficiency. Our study detailed the experimental cooling setup and demonstrated the effectiveness of the copper shield in maintaining crystal performance, a critical improvement for future solid-state nuclear clocks operating at cryogenic temperatures.
@@ 2.1 VUV photon detection setup
Our experimental setup [15, 18] was developed to produce the 229Th isomer states by using 29 keV X-ray beam and detect the signal photons, at the time when direct spectroscopy by VUV laser was unpromising [19, 20], due to the large energy uncertainty of the isomer state at that time. This report will introduce the updated setup for cooling the target crystal.
Figure 2: (a) The radioluminescence spectrum of the 229Th:CaF2 crystal measured by TU Wien group in 2015, where the distinct STE band was observed but the signal counts in Cherenkov region is barely discernible from zero [21]. The inset plot shows the measurement and calculation result for the Cherenkov region, as reported in their recent 2023 study [22]. The (a.u.) on the both vertical axes stands for arbitrary units. (b) Efficiencies of optical devices used in this measurement. Q.E.: quantum efficiency. Q.E. data were taken from Hamamatsu Photonics K. K. [23]. DCM reflectance is the total reflectance of 4 dichroic mirrors, measured by our group.
@@
To study the ice-layer formation and its impacts on isomer detection, we use the RL from the target crystal to assess the signal detection efficiency of our setup. As shown in the inset of Fig. 2 (a), the broad Cherenkov radiation spectrum includes a VUV component, which overlaps with the monochromatic isomer signal [21, 26, 22]. This VUV component can be selected using the preferential reflective band of the DCM assembly, as shown by the red curve in Fig. 2 (b). Since both the RL and isomer photons are emitted within the crystal, we use the selected Cherenkov photons as a VUV source to evaluate our setup’s efficiency for detecting isomer signals.
@@ 2.2 Cooling setup
The purpose of the cooling experiment is to lower the crystal’s temperature to approximately 100 K, reducing the damage by laser [12] or X-ray irradiation. Directly attaching the small crystal to the cold copper head or using a copper holder can achieve this target temperature; however, such methods are not feasible because we need to frequently install/uninstall, and anneal (400 C) [18] the target crystal, and we need to change the position of the crystal to gain maximum detection efficiency. Instead, we cool the stainless steel target holder to achieve the cooling of the crystal. As illustrated in Fig. 1 (c), the crystal ($\sim$1 mm3) is mounted on the front surface of the target holder, which has a hole in its body. A flexible copper wire connects the target holder to a copper rod or bar, enabling positional adjustments. The opposite end of the copper rod or bar is fixed to the cold head. The insulator and 3D stage arm are made of PEEK plastic to mitigate heat conduction from the stage.
@@ 2.3 Vacuum improvement practices
When we started the cooling experiment, the vacuum condition was 10-2 Pa, as measured with a vacuum gauge (Pfeiffer Vacuum, PKR360) at the gauge port-1 in Fig. 1 (a), and at that time the detection rate of the VUV-PMT decreased very rapidly in the initial cooling test, due to the ice-layer absorption. This relatively low vacuum condition was due to several factors: the low pumping speed of the evacuation connection, leakage of the flanges, and contamination-induced outgassing and evaporation within the chamber. This vacuum condition impedes the cooling experiments essential for the studies of 229Th isomer, necessitating efforts to address this issue.
@@ 2.4 Improved cooling scheme: copper rod conductor and shield
The shield is designed to absorb residual molecules and prevent them from condensing on the crystal surface. As shown in Fig. 1 (b), one end of the shield is closed off with the parabolic mirror with 1 mm gap kept between the mirror and the copper shield to avoid contact when the copper shrinks at low temperatures. The other end is completed with the MgF2 window which separates the vacuum between the irradiation chamber and the detection chamber. The parabolic mirror, copper shield, and MgF2 window together enclose the target holder, creating a semi-closed space free from molecular flow. Meanwhile, the cold shield works as a cold trap that captures the molecules inside it, therefore it features high vacuum conditions for the crystal cooling. The vacuum level reaches $5\times 10^{-6}$ Pa during the operation of the cooler, when the gauge was installed to the gauge port-2 in Fig. 1 (a).
@@ 3.1 Cooling test
As seen in panel (b), the introduction of the copper shield and the MgF2 window significantly suppresses the ice-layer formation, as indicated by the stable VUV-PMT counting rate when the crystal is cooled below $-100\,^{\circ}$. Therefore, during the full cooling and warming cycle, the VUV-PMT count rate was unaffected by layer absorption. The slight change in the detector counting rate near $t=3$ h could be attributed to the variation of the detection efficiency caused by multiple factors. For example, the crystal’s position could be shifted due to the contraction/expansion of the copper rod during temperature changes, and the molecules trapped in the copper shield would evaporate during the warming-up process, which presents VUV absorptions. The impact of these factors requires further studies.